# Automated Quantum Algorithm Synthesis

### v1

@@ IV.3 Performance estimation
The fitness function associates a numerical value to the performance of a candidate algorithm. The fitness is then used to generate a relative ranking of the algorithms in the population. Designing a fitness function is important as it governs how the search space is navigated. We determine the fitness of a candidate algorithm by analysing its performance with respect to set of training examples and by considering its complexity.
@@
The fitness can be improved by adding penalties to $f_{q}(m)$ that depend on the complexity of $U^{(q)}_{(m)}$. In order to favour smaller simpler structures, a penalty is added in terms of the number of unitary gates in the circuit $U^{(m)}_{i}$. A penalty is also added for the number of oracle calls and the number of gate parameters, where applicable. Lastly, we add a penalty based on the Jensen–Shannon divergence Lin (1991) which penalises circuits that perform non-uniformly across the elements of the training set. The relative weights of these penalities is a hyper-parameter of the search.
When an algorithm contains parametrised gates, these are trained numerically using gradient decentPaszke et al. (2017); Bergholm et al. (2018) with the objective of maximising the fitness of the candidate. In order to speed up evaluation, we allow parameter training to run for only a few seconds, which further penalises highly parametrised circuits. Parameters are later trained more carefully during post processing.
During evaluation the repetition parameter is increased incrementally from $1$ until the fitness no longer increases, or until the number of repetitions is equal to the circuit size. This upper-limit on the number of repetitions is motivated by the fact that we would like the size of our learnt algorithm to scale at most linearly in the number of qubits. During post-processing, the relationship between the circuit size and the number of repetitions can be inferred by evaluating the algorithm on a larger number of circuit sizes.
@@ IV.4 Tasks
In our approach, setting up the problem is the primary task of the user, while the task of designing the algorithm is delegated to the computer. As we use a bottom-up approach, the user needs to carefully construct training examples that capture the desired behaviour of the algorithm. If required, the user should also define an appropriate oracle. Here we simply used the known oracles. However, many formulations of the problem may need to be tested in order to arrive at the final algorithm, which is straightforward to do using our method. Below we discuss the problem setup for each task.
The quantum Fourier transform (QFT) Shor (1999) maps an input state to an output state whose amplitudes are related by the discrete Fourier transform. The training examples are constructed using the computational basis states and their corresponding Fourier transformed counterparts. An additional training pair is added where the input state is given by an equal superposition of all computational basis states. This penalises erroneous relative phase shifts between all other output states.
@@
Grover’s search Grover (1997) provides an optimal method for searching an unstructured database. The oracle marks the state associated with the sought database entry with a phase, namely $\exp(i\pi)=-1$. A training pair is constructed for each of the possible sought entries (computational basis states) for the given problem size $2^{q}$, where $q$ is the number of qubits. A fixed input state of (|$(\frac{\ket{0}+\ket{1}}{\sqrt{2}})^{\otimes q}$ is used, which encodes the assumption that each database entry has a equal probability being the sought entry.
@@ V Data Availability
The dataset generated and analysed during the current study is not publicly available due to its size but is available from the corresponding author on reasonable request.

### v2

@@ IV.3 Performance estimation
The fitness function associates a numerical value to the performance of a candidate algorithm. The fitness is then used to generate a relative ranking of the algorithms in the population. Designing a fitness function is important as it governs how the search space is navigated. We determine the fitness of a candidate algorithm by analysing its performance with respect to a set of training examples and by considering its complexity.
@@
The fitness can be improved by adding penalties to $f_{q}(m)$ that depend on the complexity of $U^{(q)}_{(m)}$. To favour smaller simpler structures, a penalty is added in terms of the number of unitary gates in the circuit $U^{(m)}_{i}$. A penalty is also added for the number of oracle calls and the number of gate parameters, where applicable. Lastly, we add a penalty based on the Jensen–Shannon divergence Lin (1991) which penalises circuits that perform non-uniformly across the elements of the training set. The relative weights of these penalties are a hyper-parameter of the search.
When an algorithm contains parametrised gates, these are trained numerically using gradient decent  Paszke et al. (2017); Bergholm et al. (2018) to maximise the fitness of the candidate. In order to speed up evaluation, we allow parameter training to run for only a few seconds, which further penalises highly parametrised circuits. Parameters are later trained more carefully during post-processing.
During evaluation, the repetition parameter is increased incrementally from $1$ until the fitness no longer increases, or until the number of repetitions is equal to the circuit size. This upper limit on the number of repetitions is motivated by the fact that we would like the size of our learnt algorithm to scale at most linearly in the number of qubits. During post-processing, the relationship between the circuit size and the number of repetitions can be inferred by evaluating the algorithm on a larger number of circuit sizes.
@@ IV.4 Tasks
In our approach, setting up the problem is the primary task of the user, while the task of designing the algorithm is delegated to the computer. As we use a bottom-up approach, the user needs to carefully construct training examples that capture the desired behaviour of the algorithm. If required, the user should also define an appropriate oracle. Here we simply used the known oracles. However, many formulations of the problem may need to be tested to arrive at the final algorithm, which is straightforward to do using our method. Below we discuss the problem setup for each task.
The quantum Fourier transform (QFT) Shor (1999) maps an input state to an output state whose amplitudes are related by the discrete Fourier transform. The training examples are constructed using the computational basis states and their corresponding Fourier-transformed counterparts. An additional training pair is added where the input state is given by an equal superposition of all computational basis states. This penalises erroneous relative phase shifts between all other output states.
@@
Grover’s search Grover (1997) provides an optimal method for searching an unstructured database. The oracle marks the state associated with the sought database entry with a phase, namely $\exp(i\pi)=-1$. A training pair is constructed for each of the possible sought entries (computational basis states) for the given problem size $2^{q}$, where $q$ is the number of qubits. A fixed input state of (|$(\frac{\ket{0}+\ket{1}}{\sqrt{2}})^{\otimes q}$ is used, which encodes the assumption that each database entry has an equal probability being the sought entry.
@@ V Data Availability
The dataset generated and analysed during the current study is not publicly available due to its size but is available from the corresponding author upon reasonable request.